Write each equation in Slope-Intercept Form if possible and identify the slope and y-intercept if they exist.
6x-4y=24
The answer I obtained was y=-6/4x + 24/4
y-int: 24/-4

If I got this incorrect can a tutor please explain what step I did incorrectly?? Thank you!

Answer by Boreal(15235)

Close. x=0, y=-6 (0, -6) and y=0 x=4 (4, 0)
It is -4y=-(6/4)x+24
y=(3/2)x-6
x and y will both have the same sign on the opposite side of the equal sign.
The constant will be negative for a positive y.